OpenStudy (anonymous):

Find the value of y. -6y + 14 + 4y = 32

OpenStudy (bradely):

-6y+14+4y=32 First arrange the like terms

OpenStudy (anonymous):

y= -9 because combine like terms -6y+4y= -2y then u subtract 14 from both sides 32-14 = 18 then u divide by -2 on both sides 18/-2 y-9

OpenStudy (bradely):

-6y+4y+14=32 Using the distributive law (-6+4)y+14=32

OpenStudy (bradely):

-2y+14=32 Subtract 14 both sides -2y=32-14 -2y=18 y=-9 any help http://www.mathskey.com/question2answer/ask